OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 21 vowels and 15 consonants, what is the probability you will choose a consonant and then a vowel?

OpenStudy (anonymous):

@amistre64 help :)

OpenStudy (anonymous):

1/36?

OpenStudy (anonymous):

Total outcomes = 36 Favorable outcomes = 15 probability you will choose a consonant = Total outcomes/Favorable outcomes = 15/36 = 5/12 Total outcomes = 36 Favorable outcomes = 21 probability you will choose a vowel = 21/36 = 7/12 Probability you will choose a consonant and then a vowel = 5/12 * 7/12 = 35/144